Suppose the functions f,g and h are defined by f(x)=3^x+3^-x+3 and g(x)=log5(x+2)+log5(3)-log4(2x-3) and h(x)=e^x+15e^-x respectively 4.1 write down the sets Df,DG and Dh that represent the domains of f g and h respectively. 4.2 solve the equation f(x)=12 for x. 4.3 solve the inequality g(x)= 0 for x. 4.4solve h(x)=8 for x. Leave the answer in terms of ln, where necessary


1
Expert's answer
2022-06-13T15:32:24-0400

4.1


"f(x)=3^x+3^{-x}+3"

"Df: (-\infin, \infin)"



"g(x)=\log_4(x+2)+\log_4(3)-\log_4(5)-\log_4(2x-3)""x+2>0=>x>-2""2x-3>0=>x>\dfrac{3}{2}"

"Dg:(\dfrac{3}{2}, \infin)"




"h(x)=e^x+15e^{-x}"

"Dh: (-\infin, \infin)"


4.2


"f(x)=12"

"3^x+3^{-x}+3=12"

"(3^x)^2-9(3^x)+1=0"

"Determinant=(-9)^2-4(1)(1)=77"

"3^x=\dfrac{9\pm\sqrt{77}}{2(1)}"

"x_1=\dfrac{\ln(\dfrac{9-\sqrt{77}}{2})}{\ln3},"

"x_2=\dfrac{\ln(\dfrac{9+\sqrt{77}}{2})}{\ln3}"

4.3


"\log_4(x+2)+\log_4(3)-\log_4(5)-\log_4(2x-3)=0"

"\log_4\dfrac{3(x+2)}{5(2x-3)}=0"

"\dfrac{3(x+2)}{5(2x-3)}=1"

"3x+6=10x-15"

"x=3"

4.4


"e^x+15e^{-x}=8"

"(e^x)^2-8(e^x)+15=0"

"(e^x-3)(e^x-5)=0"

"e^x-3=0\ or\ e^x-5=0"

"x_1=\ln 3, x_2=\ln 5"